Four typists in a room guess rise to an average intensity level of 50db. What is the intensity level in the room when an additional two men and two women generate the same amount of noise.
1
Expert's answer
2018-02-05T08:16:08-0500
L=50 dB=10 log⁡〖4I/I_0 〗
L^'=10 log⁡〖8I/I_0 〗=10(log⁡〖4I/I_0 〗+log⁡2 )=10(log⁡〖4I/I_0 〗+log⁡2 )=50+10 log⁡2=53 dB.